### v1

# Spatial Proportional Hazards Model with Differential Regularization

Spatial Proportional Hazards Model with Differential Regularization
Lorenzo Tedesco
1 Introduction
Survival analysis has long been an essential tool across various fields, including medicine, engineering, and social sciences, for analyzing time-to-event data. In this paper, we focus on one of the most widely used models within survival analysis: the Proportional Hazards model, first introduced by Cox (1972). This model assesses the effect of covariates on the hazard, or risk, of an event occurring.
In recent years, survival analysis has made significant strides in modeling different forms of dependence in survival data. This paper explores a scenario where the occurrence of an event is associated with its geographic location, introducing a spatial dependence into the data. Specifically, an event’s occurrence may be influenced by spatial factors related to the location where the event happens. We refer to the Proportional Hazards model that incorporates these spatial factors as the Spatial Proportional Hazards model.
Spatial factors can generally be categorized into two main approaches: the geostatistical approach, which uses geographic coordinates (latitude and longitude), and the lattice approach, which considers the relative position of an area to neighboring regions and our method is able to deal with both the framework. A comprehensive literature review, including Bayesian approaches, can be found in Hanson and Zhou (2014). Our interest lies in a frequentist approach to the spatial Proportional Hazards model, for which various models and methods are now available to facilitate inference. Examples include the semiparametric estimating equations method proposed by Li and Lin (2006) and the composite likelihood method Paik and Ying (2012). The main issue with these proposals is that the dependence structure does not consider the shape of the domain but only the distance among the locations. This can be problematic when the boundary of the domain plays a crucial role in the spatial dependence. In this paper, we propose a novel approach based on recent developments in differential regularization, as discussed in Wilhelm and Sangalli (2016). One of the main features of this approach is to be able to deal with domains having complex shapes, characterized for instance by strong concavities or holes, and where the shape of the domain influences the behavior of the phenomenon.
We demonstrate that the methodology introduced in Wilhelm and Sangalli (2016), which concerns a generalized regression model for spatially distributed data when the response variable has a distribution within the exponential family, can be extended to a spatial proportional hazard model with randomly right-censored data. Specifically, we optimize a penalized log-likelihood function that includes a roughness penalty term based on a differential quantity of the spatial field, calculated over the domain of interest. This approach utilizes the so called Generalized Spatial Regression with PDE penalization (GSR-PDE) method (Wilhelm and Sangalli, 2016). Our proposed models leverage finite element methods on a triangulation of the domain to achieve precise estimates of the spatial field, particularly in domains with complex geometries, where triangulation effectively captures the intricacies of the area.
The paper is organized as follows: In Section 2, we introduce the Spatial Proportional Hazards model. Section 3 discusses the functional version of the PIRLS, tailored to our specific case of interest. In Section 4, we describe the numerical implementation of the method using the finite element method. Section 5 specifies the model version adapted for areal data. Section 6 is dedicated to simulation studies, and Section 7 presents an empirical application of the method. Finally, Section 8 provides the conclusions. All technical details and proofs are deferred to the Appendix.
2 Model
We consider a bounded domain $\Omega\subset\mathbb{R}^{2}$ with a boundary $\partial\Omega$ that is regular and belongs to $C^{2}(\mathbb{R}^{2})$. Let $p_{1},\ldots,p_{n}\in\Omega$ denote $n$ points, where $p_{i}=(p_{1i},p_{2i})$. Each point $p_{i}$ is associated with a time-to-event random variable $T_{i}$. Due to censoring, $T_{i}$ cannot be observed directly; instead, we obtain a realization of $(Y_{i},\Delta_{i})$, denoted by $(y_{i},\delta_{i})$, where $Y_{i}=\min(T_{i},C_{i})$, $\Delta_{i}=I(T_{i}\leq C_{i})$, $C_{i}$ is a censoring random variable, and $I(\cdot)$ is the indicator function. Additionally, we observe a covariate vector $x_{i}\in\mathbb{R}^{q}$ at each location $p_{i}$. We assume that $T_{i}$ is independent of $C_{i}$ conditional on $x_{i}$. We use a proportional hazard model for the random variables $\{T_{i}\}_{i=1}^{n}$. Specifically, let $\lambda_{i}(t\mid x)$ denote the conditional hazard function of $T_{i}$ given $x_{i}=x$. We assume that
$\lambda_{i}(t\mid x)=\lambda_{0}(t)\exp(x^{\top}\beta+f(p_{i})),$ (1)
where $\lambda_{0}$ is the unspecified baseline hazard, $\beta\in\mathbb{R}^{q}$ is a vector of regression coefficients, and $f$ is a smooth function over $\Omega$ belonging to a suitable functional space $\mathcal{F}$. Note that $f$ is identifiable up to an additive constant, so we assume that $\mathcal{F}$ contains only functions such that at a given point $p^{*}\in\Omega$ it holds $f(p^{*})=0$.
We propose to estimate the regression coefficients $\beta\in\mathbb{R}^{q}$ and the spatial field $f\in\mathcal{F}$ by maximizing a penalized log-partial-likelihood functional:
[…]
where $\mathcal{L}(\beta,f)$ is the standard partial likelihood defined as
$\mathcal{L}(\beta,f)=\sum_{i:\delta_{i}=1}\left(x_{i}^{\top}\beta+f(p_{i})-\log\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\beta+f(p_{j})\right)\right).$ (3)
Here, $\mu$ is a positive smoothing parameter, and the Laplacian $\Delta f=\frac{\partial^{2}f}{\partial p_{1}^{2}}+\frac{\partial^{2}f}{\partial p_{2}^{2}}$ measures the local curvature of the field $f$. A higher $\mu$ controls the wiggliness of $f$, while a smaller $\mu$ allows for more flexibility in $f$.
We show that maximizing the penalized log-likelihood function (2) is equivalent to minimizing the penalized least-squares functional considered in Sangalli et al. (2013). The quadratic form of this functional allows for an analytical characterization of its minimum (or equivalently, the maximum of the penalized log-likelihood functional), and thus for the estimation of $\hat{\beta}\in\mathbb{R}^{q}$ and $\hat{f}\in\mathcal{F}$. As in Wilhelm and Sangalli (2016), which considers an exponential family, we cannot directly apply the standard PIRLS algorithm developed by O’sullivan et al. (1986) for maximizing a penalized log-likelihood functional in generalized additive models. This is because the penalized log-likelihood in (2) involves a function parameter, the spatial field $f$, and the maximization occurs over the space $\mathbb{R}^{q}\times\mathcal{F}$, where $\mathcal{F}$ is infinite-dimensional. Instead we recall the functional version of the PIRLS algorithm proposed in Wilhelm and Sangalli (2016) to find an approximate solution to the estimation problem.
3 Functional version of the PIRLS algorithm
We consider the more general penalized log-likelihood function given by:
$\mathcal{L}_{p}(\beta,f)=\mathcal{L}(\beta,f)-\frac{\mu}{2}m(f,f),$ (4)
where $\mathcal{L}$ denotes the log-likelihood and $m(\cdot,\cdot):\mathcal{F}\times\mathcal{F}\to\mathbb{R}$ is a bilinear, symmetric, and semi-positive definite form. This formulation enables us to extend the PIRLS algorithm to accommodate any functional roughness penalty of this general quadratic form.
For that, define the following quantities:
$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
$r_{ji}$ $=I(Y_{j}\geq Y_{i}),$
$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f),$
$l_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)},$
$v_{i}(\beta,f)$ $=\theta_{i}(\beta,f)/l_{i}(\beta,f).$
First, we show that the problem of maximizing 4 is equivalent to minimizing the following functional $\mathcal{J}_{\mu}(\beta,f)$ with respect to $(\beta,f)$:
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$, where such quantities are supposed to be constant. As, in fact, $l$ and $V$ also depend on $(\beta,f)$, this suggests an iterative scheme for the solution of the estimation problem. Let $(\beta^{(k)},f_{n}^{(k)})$ be the value of $(\beta,f)$ after $k$ iterations of the algorithm, and let $\theta^{(k)}$ the estimate of the vector $\theta$ after $k$ iterations, and define $\Theta^{(k)}=\text{diag}(\theta^{(k)})$. Define also $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X^{\top}\beta^{(k)}+f^{(k)}$, and let ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$. The first order approximation of $\theta(\beta,f)$ in the neighborhood of the current value of $\theta^{(k)}$ has to be considered in the space $\mathbb{R}^{q}\times\mathcal{F}$ and yields to the following quadratic approximation of $\mathcal{J}_{\mu}(\beta,f)$:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X^{\top}\beta-f)\right\|^{2}+\mu m(f,f).$ (5)
We can consider the following iterative scheme:
  1. 1.
    Compute $z^{(k)}$ and $W^{(k)}$ based on $(\beta^{(k)},f_{n}^{(k)})$.
  2. 2.
    Next, find $\beta$ and $f_{n}$ that minimize (5).
  3. 3.
    Perform the standard proportinal hazards model estimation using $X$ and $f_{n}^{(k)}$ as covariate and denote by $\hat{\beta}$ and $\hat{c}$ the resulting regression coeffiecint for $X$ and $f_{n}^{(k)}$, respectively.
  4. 4.
    Define $\beta^{(k+1)}=\hat{\beta}$ and and $f_{n}^{(k+1)}=\hat{c}(f_{n}^{(k)}-f_{n}^{(k)}(p^{*}))$.
Here $f_{n}^{(k)}(p^{*})$ is the estimation of $f$ at step $k$ at $p^{*}$. The standardization of $f_{n}^{(k)}$ in step 4 ensures that $f_{n}(p^{*})=0$ and that (1) is satisfied. We can use a stopping criteria based on a sufficiently small variation of two successive values of (5) or a maximum number of iterations.
4 Numerical Implementation using the Finite Element Method
[…]
$\tilde{J}_{\lambda}(\beta,f)=\|W^{1/2}(z-X\beta-f_{n})\|^{2}+\lambda\int_{\Omega}(\Delta f)^{2}.$ (6)
For simplicity, we omit the dependence on the iteration counter $k$ in this notation and throughout the following sections. Let us now consider what type of space $\mathcal{F}$ is suitable for the current problem. To do so, we introduce the Sobolev space $H^{m}(\Omega)$, which is a Hilbert space comprising all functions in $L^{2}(\Omega)$ along with all their distributional derivatives up to order $m$. Since the roughness penalty term $\int_{\Omega}(\Delta f)^{2}$ must be well-defined, it follows that $\mathcal{F}\subset H^{2}(\Omega)$. According to the Sobolev embedding theorem, $H^{2}(\Omega)\subset C^{0}(\Omega)$. Therefore, any function $f\in H^{2}(\Omega)$ is continuous and can be evaluated pointwise, allowing us to compute the vector $f_{n}$ in the partial likelihood term. To ensure the uniqueness of the minimizer in (6), appropriate boundary conditions must be imposed. These conditions specify the desired behavior of the estimated function $f$ at the boundaries of the domain of interest. Typically, constraints can be set on the value of $f$ at the boundary $\partial\Omega$, i.e., $f|_{\partial\Omega}=\gamma_{D}$ (Dirichlet boundary conditions), or on the flux of the function through the boundary, i.e., $\partial_{n}f|_{\partial\Omega}=(\nabla f)^{\top}\mathbf{n}=\gamma_{N}$ (Neumann boundary conditions), where $\mathbf{n}$ is the outward-pointing normal unit vector to the boundary, and $\nabla f=(\partial f/\partial p_{1},\partial f/\partial p_{2})^{\top}$ represents the gradient of $f$. If $\gamma_{D}$ or $\gamma_{N}$ are zero functions, the boundary conditions are termed homogeneous. It is also possible to apply different boundary conditions to various parts of the boundary, forming a partition of $\partial\Omega$. To ensure the uniqueness of the minimization problem (4.1), we consider the space:
$\mathcal{F}=H^{2}_{n}=\left\{f\in H^{2}\mid(\nabla f)^{\top}\mathbf{n}=0\text{on }\partial\Omega\right\}.$
For more details on general boundary conditions, see Azzimonti et al. (2014).
4.1 Characterization of the Solution to the Penalized Partial Likelihood Problem
We assume that the design matrix $XX^{\top}$ is of full rank and that the weight matrix $W$ contains strictly positive entries. Define $H=X(X^{\top}WX)^{-1}X^{\top}W$ and $Q=I-H$, where $I$ is an identity matrix of the appropriate size. For any function $u$ within the functional space $\mathcal{F}=H^{2}_{n}(\Omega)$, let $u_{n}=(u(p_{1}),\ldots,u(p_{n}))$ denote the vector of evaluations of $u$ at the $n$ spatial locations $\{p_{i}\}_{i=1}^{n}$. Let $\tilde{\beta}$ and $\tilde{f}$ denote the minimizers of the functional $\tilde{J}_{\lambda}(\beta,f)$ in (6), and $\hat{\beta}$ and $\hat{f}$ the maximizers of the penalized partial likelihood functional $\mathcal{L}_{p}(\beta,f)$ in (4). Under these conditions, the following Proposition describes the minimizers $\tilde{\beta}$ and $\tilde{f}$ of the functional (6).
Proposition 1.
There exists a unique pair $(\tilde{\beta},\tilde{f})\in\mathbb{R}^{q}\times H^{2}_{n}$ that minimizes (4). Furthermore,
  • $\tilde{\beta}=(X^{\top}WX)^{-1}X^{\top}W(z-\tilde{f}_{n})$, where $\tilde{f}_{n}=(\tilde{f}(p_{1}),\ldots,\tilde{f}(p_{n}))^{\top}$,
  • $\tilde{f}$ satisfies:
    $u_{n}^{\top}Q\tilde{f}_{n}+\lambda\int_{\Omega}(\Delta u)(\Delta\tilde{f})=u_{n}^{\top}Qz,\quad\forall u\in H^{2}_{n}.$ (7)
Using Proposition 1 and the functional version of the PIRLS algorithm from Section 3, we obtain a characterization of the maximum penalized log-likelihood within the functional space $H^{2}_{n}(\Omega)$.
4.2 Solution to the Penalized Partial Likelihood
Here, we outline the methodology for minimizing $\tilde{J}_{\lambda}(\beta,f)$ with respect to both $\beta$ and $f$. As indicated by Proposition 1, computing $\tilde{\beta}$ is straightforward. The key challenge is to find $\tilde{f}$ that satisfies (7). To address this, we introduce the space:
$H^{1}_{n}=\left\{f\in H^{1}\mid(\nabla f)^{\top}\mathbf{n}=0\text{ on }\partial\Omega\right\}.$
According to Wilhelm and Sangalli (2016); Sangalli et al. (2013), problem (7) is equivalent to finding $(\tilde{f},\tilde{h})\in H^{1}_{n}(\Omega)\times H^{1}_{n}(\Omega)$ such that
$\begin{cases}u_{n}^{\top}Q\tilde{f}_{n}-\lambda\int_{\Omega}(\nabla u)^{\top}\nabla\tilde{h}=u_{n}^{\top}Qz,\\ -\int_{\Omega}(\nabla\tilde{f})^{\top}\nabla v=\int_{\Omega}\tilde{h}v,\end{cases}$ (8)
for each $(u,v)\in H^{1}_{n}\times H^{1}_{n}(\Omega)$. This formulation requires less regularity on the functions involved compared to formulation (6) defined in $H^{2}_{n}(\Omega)$. In the following section, we discuss how the finite element method can be employed to construct a finite-dimensional subspace of $H^{1}_{n}(\Omega)$, enabling the computation of an approximate solution to (7) within this space.
4.3 Finite Elements
The finite element method is a prevalent numerical technique in engineering for solving problems involving partial differential equations (Quarteroni and Quarteroni, 2009).
To build a finite element space, we begin by dividing the domain of interest $\Omega$ into smaller subdomains. Common domain partitions include triangular meshes. We consider a regular triangulation $\mathcal{T}$ of $\Omega$, where adjacent triangles share either a vertex or an edge. The domain $\Omega$ is thus approximated by $\Omega_{\mathcal{T}}$, which is the union of all triangles. Consequently, the boundary $\partial\Omega$ is approximated by a polygon (or multiple polygons for domains with interior holes). This triangulation effectively captures the complex geometry of the domain, including significant concavities such as rivers and detailed structures like bridges that connect different parts of the city.
From this triangulation, locally supported polynomial functions are defined over the triangles, forming a set of basis functions $\psi_{1},\ldots,\psi_{K}$ that span a finite-dimensional subspace $\mathcal{F}_{K}$ of $H^{1}_{n}$. Linear finite elements are created using a basis system where each basis function $\psi_{i}$ is associated with a vertex $\xi_{i}$ (for $i=1,\ldots,K$) of the triangulation $\mathcal{T}$. Each basis function $\psi_{i}$ is a piecewise linear polynomial that equals one at the vertex $\xi_{i}$ and zero at all other vertices of the mesh, i.e., $\psi_{i}(\xi_{j})=\delta_{ij}$ for all $i,j=1,\ldots,K$, where $\delta_{ij}$ is the Kronecker delta. Figure 1 provides an example of a linear finite element basis function on a planar mesh, highlighting its localized support.
Let $\psi=(\psi_{1},\ldots,\psi_{K})^{\top}$ be the column vector containing the $K$ piecewise linear basis functions corresponding to the $K$ vertices $\xi_{1},\ldots,\xi_{K}$. Any function $h$ in the finite element space $\mathcal{F}_{K}$ can be expressed as a linear combination of the basis functions $\psi_{1},\ldots,\psi_{K}$. Let $h=(h_{1},\ldots,h_{K})$ represent the coefficients of this basis expansion for $h$, such that
$h(\cdot)=\sum_{j=1}^{K}h_{j}\psi_{j}(\cdot)=\mathbf{h}^{\top}\psi(\cdot).$
Note that
$h(\xi_{i})=\sum_{j=1}^{K}h_{j}\psi_{j}(\xi_{i})=\sum_{j=1}^{K}h_{j}\delta_{ij}=h_{i},$
therefore,
$\mathbf{h}=(h(\xi_{1}),\ldots,h(\xi_{K})),$
which shows that each function $h\in\mathcal{F}_{K}$ is completely determined by its values at the mesh nodes.
Figure 1: Example of a linear finite element basis function on a planar mesh. The function is locally supported, taking the value one at a specific vertex and zero at all other vertices of the mesh.
4.4 Numerical solution to the Penalized Partial Likelihood
The functions and integrals in equation (8) can be approximated using elements from the finite element space $\mathcal{F}_{K}$. As a result, the problem (8) is approximated by its discrete version: find $(\tilde{f},\tilde{h})\in\mathcal{F}_{K}\times\mathcal{F}_{K}$ such that (8) holds for all $(u,v)\in\mathcal{F}_{K}\times\mathcal{F}_{K}$, where the integrals are now computed over the triangulation $\Omega_{\mathcal{T}}$. Define $\Psi$ as the $n\times K$ matrix representing the evaluations of the $K$ basis functions at the $n$ data locations $\mathbf{p}_{1},\ldots,\mathbf{p}_{n}$,
$\Psi=\begin{bmatrix}\psi^{\top}(\mathbf{p}_{1})\\ \vdots\\ \psi^{\top}(\mathbf{p}_{n})\end{bmatrix}$ (9)
and consider the $K\times K$ matrices
$\mathbf{R}_{0}=\int_{\Omega_{\mathcal{T}}}(\psi\psi^{\top}),\quad\mathbf{R}_{1}=\int_{\Omega_{\mathcal{T}}}(\nabla\psi)^{\top}\nabla\psi.$
Using this notation, for functions $\tilde{f},\tilde{h},u,v\in\mathcal{F}_{K}$, the integrals in (8) can be written as:
$\int_{\Omega_{\mathcal{T}}}(\nabla u)^{\top}\nabla\tilde{h}=\mathbf{u}^{\top}\mathbf{R}_{1}\tilde{\mathbf{h}},\quad\int_{\Omega_{\mathcal{T}}}(\nabla\tilde{f})^{\top}\nabla v=\tilde{\mathbf{f}}^{\top}\mathbf{R}_{1}\mathbf{v},\quad\int_{\Omega_{\mathcal{T}}}\tilde{h}\,v=\tilde{\mathbf{h}}^{\top}\mathbf{R}_{0}\mathbf{v},$
where $\tilde{\mathbf{f}},\tilde{\mathbf{h}},\mathbf{u}$, and $\mathbf{v}$ are the coefficient vectors of the basis expansions of $\tilde{f},\tilde{h},u$, and $v$, respectively. Consequently, the discrete version of problem (8) reduces to solving a linear system, as formulated in the following proposition.
Proposition 2.
The discrete form of (8) is given by the system
$\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\Psi&\lambda\mathbf{R}_{1}\\ \lambda\mathbf{R}_{1}&\lambda\mathbf{R}_{0}\end{bmatrix}\begin{bmatrix}\tilde{\mathbf{f}}\\ \tilde{\mathbf{h}}\end{bmatrix}=\begin{bmatrix}-\Psi^{\top}\mathbf{Q}\mathbf{z}\\ 0\end{bmatrix},$ (10)
which has a unique solution pair $(\tilde{\mathbf{f}},\tilde{\mathbf{h}})$ corresponding to the coefficients of the basis expansions of $\tilde{f}$ and $\tilde{h}$.
The uniqueness of the solution to (10) is guaranteed by the positive definiteness of the matrices $\mathbf{R}_{0}$ and $(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1})$.
Let $\mathbf{P}=\mathbf{R}_{1}\mathbf{R}_{0}^{-1}\mathbf{R}_{1}$ and $\mathbf{S}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}$. Using the functional form of the PIRLS algorithm, and leveraging Propositions 1 and 2, we obtain the following expressions for the maximizers $\hat{\bm{\beta}}$ and $\hat{f}$ of the penalized partial likelihood (2):
$\hat{\bm{\beta}}=(\mathbf{X}^{\top}\mathbf{W}\mathbf{X})^{-1}\mathbf{X}^{\top}\mathbf{W}(\mathbf{I}-\mathbf{S})\mathbf{z},$
with $\hat{f}$ identified by the vector
$\hat{\mathbf{f}}=(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z},$
and the vector of evaluations of $\hat{f}$ at the $n$ data locations given by
$\hat{\mathbf{f}}_{n}=\Psi\hat{\mathbf{f}}=\Psi(\Psi^{\top}\mathbf{Q}\Psi+\lambda\mathbf{P})^{-1}\Psi^{\top}\mathbf{Q}\mathbf{z}=\mathbf{S}\mathbf{z},$
where the vector of pseudo-data $\mathbf{z}$, and the matrices $\mathbf{W},\mathbf{Q}$, and $\mathbf{S}$ are those obtained upon convergence of the PIRLS algorithm.
The positive definite matrix $\mathbf{P}$ represents the discretization of the penalty term in (2) and (7). Due to the variational formulation (8) of the estimation problem, this penalty matrix does not require the computation of second-order derivatives. Azzimonti et al. (2014) demonstrates that in the finite element space used for discretization, $\mathbf{P}$ is equivalent to the penalty matrix that would result from a direct discretization of the penalty term in (2) and (7), which would otherwise involve the computation of second-order derivatives.
4.5 Selecting the Parameter $\mu$ Using Cross-Validation
When it comes to choosing the parameter $\mu$, one straightforward approach is a visual inspection. This method allows the researcher to select a value that achieves the desired smoothness of the estimated spatial field. However, to avoid subjective decision-making, we propose a more objective, numerical method. Following the approach of Cygu et al. (2021), we can utilize cross-validation to determine the optimal value of $\mu$. Specifically, we select the $\mu$ that results in the lowest cross-validated partial likelihood deviance (CV-PLD) or the highest cross-validated Harrell’s concordance index (CV-C-index) (Harrell Jr et al., 1996).
To compute the CV-PLD for each $\mu$, we employ $k$-fold cross-validation. This process involves splitting the training data into $k$ subsets (folds), training the model on $k-1$ folds, and validating it on the remaining fold. This procedure is repeated $k$ times, each time with a different fold serving as the validation set. We use two metrics for this purpose: CV-PLD and CV-C-index (Dai and Breheny, 2019). The CV-PLD is defined as:
[…]
where $\mathcal{L}(\hat{\beta}_{-k}(\mu))$ represents the log partial likelihood evaluated at $\hat{\beta}_{-k}$ using the entire dataset, and $\mathcal{L}_{-k}(\hat{\beta}_{-k}(\mu))$ is the log partial likelihood evaluated at $\hat{\beta}_{-k}$ on the retained data (all but the excluded fold). The $\hat{\beta}_{-k}$ values denote the penalized estimates derived from the retained data. The optimal value of $\mu$ is the one that minimizes (11). Notably, equation (11) typically provides superior results compared to simply evaluating the partial likelihood on the held-out fold (a method sometimes referred to as the basic approach), as the likelihood of any given observation is influenced by the other elements in the risk set.
The alternative cross-validation metric, the CV-C-index, is based on Harrell’s concordance index for Cox models. This metric, known as the cross-validated C-index, calculates the probability that, for a random pair of individuals, the predicted survival times are correctly ordered according to their actual survival times. This method is similar to the implementation available in the survival package (Therneau et al., 2015).
5 Model Version for Areal data
//TODO
6 Simulation Results
To demonstrate the performance of the proposed model, we present simulations on a horseshoe domain, as defined in Ramsay (2002); Wood et al. (2008). We use the spatial field $f(x,y)$, depicted in Figure 2(a) and defined as follows:
$f(x,y)=\begin{cases}-\frac{1}{8}\left(x+\frac{\pi r}{2}+(y-r)^{2}+10\right)+c&\text{if }(x-3)^{2}+(y-r)^{2}\leq(r-r_{0})^{2}\text{ and }x\geq 3,\\ -\frac{1}{8}\left(x+\frac{\pi r}{2}+(y-r)^{2}+10\right)+c&\text{if }0\leq x\leq 3,\,r_{0}\leq y\leq 2r-r_{0},\\ -\frac{1}{8}\left(\arctan\left(-\frac{y}{x}\right)r+\left(\sqrt{x^{2}+y^{2}}-r\right)^{2}+10\right)+c&\text{if }r_{0}^{2}\leq x^{2}+y^{2}\leq(2r-r_{0})^{2}\text{ and }x<0,\\ -\frac{1}{8}\left(-x-\frac{\pi r}{2}+(-y-r)^{2}+10\right)+c&\text{if }0\leq x\leq 3,\,-(2r-r_{0})\leq y\leq-r_{0},\\ -\frac{1}{8}\left(-x-\frac{\pi r}{2}+(-y-r)^{2}+10\right)+c&\text{if }(x-3)^{2}+(y+r)^{2}\leq(r-r_{0})^{2}\text{ and }x>3,\\ 0&\text{otherwise},\end{cases}$
where $c$ is a constant chosen so that the minimum value of $f$ over the horseshoe domain is zero, approximately $c\simeq 1.77$, $r=0.5$ and $r_{0}=0.1$.
We randomly sample $n=100,200,300$ points, denoted by $\{p_{i}\}_{i=1}^{n}$, from the horseshoe domain, as illustrated in Figure 2(b). Figure 2(c) shows the corresponding Delaunay triangulation. Each sampled point is associated with an observation $x_{i}$ of a random variable $X_{i}$, which follows a standard normal distribution. Additionally, we observe $(y_{i},\delta_{i})$, where $y_{i}=\min(T_{i},C_{i})$ and $\delta_{i}=I(T_{i}\leq C_{i})$. The variable $T_{i}$ given $X_{i}=x$ follows a proportional hazards model with baseline hazard $\lambda_{0}(t)=1$ and $\beta_{0}=0.2$, as defined in equation (1), with the form $\lambda(t|x)=\lambda_{0}(t)\exp(x\beta_{0}+f(p_{i}))$. To introduce censoring, we set $C_{i}=1$ for each $i$, resulting in approximately 20% censored observations. We set $\mu=0.5,1,5,10$ in order to understand how the value of $\mu$ effects the estimation of the spatial field $f$ and of the regression coefficients vector of $\beta$. The finite elements method is applied using a basis of order $1$ constructed on the Toggle the table of contents Delaunay triangulation over the sampled locations.
(a) Spatial field $f(x,y)$ over the horseshoe domain
(b) Sampled points over the horseshoe domain
(c) Mesh based on the sampled points
Figure 2: Figures illustrating various aspects of the spatial field and sampling.
We conduct $N=100$ simulations for each combination of $n$ and $\mu$. And example of the estimated spatial field in Figure 3. This figure compares the true values (Figure 3(a)) with the estimated values (Figure 3(b)), and reports the pointwise error. In table
(a) True Values
(b) Estimated Values
(c) Error Values
Figure 3: Comparison of true values, estimated values, and error values.
$n$ $\mu$ MSE of $f$ (sd) Bias of $\beta$ (sd)
100 0.5 0.168 (0.112) -0.008 (0.130)
1 0.168 (0.118) -0.007 (0.130)
3 0.177 (0.140) -0.006 (0.130)
5 0.186 (0.165) -0.006 (0.130)
10 0.190 (0.188) -0.007 (0.129)
200 0.5 0.157 (0.119) 0.008 (0.096)
1 0.160 (0.129) 0.007 (0.096)
3 0.168 (0.150) 0.007 (0.095)
5 0.180 (0.169) 0.007 (0.094)
10 0.185 (0.179) 0.008 (0.093)
300 0.5 0.141 (0.105) -0.003 (0.077)
1 0.142 (0.115) -0.002 (0.077)
3 0.143 (0.129) 0.000 (0.077)
5 0.144 (0.136) 0.000 (0.077)
10 0.148 (0.149) 0.000 (0.077)
Table 1: Simulation results for varying sample sizes ($n$) and smoothing parameters ($\mu$), showing the mean squared error (MSE) of the spatial field estimator $f$ and the bias of the parameter $\beta$. Standard deviations (sd) are reported in parentheses for both MSE and bias.
7 Empirical Application
//TODO
8 Conclusions
//TODO
Appendix A Functional PIRLS Algorithm details
The maximizer $(\hat{\beta},\hat{f})$ of (4) must satisfy the following system of first-order equations:
$\begin{cases}\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]-\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (12)
This system involves derivatives with respect to both parameters $\beta$ and $f$, where the derivative with respect to $f$ is the Gâteaux derivative in the direction of $u\in\mathcal{F}$.
For each $k=1,\dots,q$, the term $\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}$ is given by:
$\frac{\partial\mathcal{L}(\hat{\beta},\hat{f})}{\partial\beta_{k}}=\sum_{i:\delta_{i}=1}\left(x_{ik}-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)x_{jk}}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (13)
For the second equation of (12), we note that $m(u,\hat{f})$ is the derivative of $m(\hat{f},\hat{f})$. We now compute the terms involving $\mathcal{L}$ only:
$\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})$ $=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{i})+tu(p_{i})\right)}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)(1+tu(p_{i})+o(t))}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-\log\left\{1+t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right\}\right)$
$=\sum_{i:\delta_{i}=1}\left(tu(p_{i})-t\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}+o(t)\right).$
Thus, we obtain:
$\lim_{\tau\to 0}\frac{1}{\tau}\left[\mathcal{L}(\hat{\beta},\hat{f}+\tau u)-\mathcal{L}(\hat{\beta},\hat{f})\right]$ $=\sum_{i:\delta_{i}=1}\left(u(p_{i})-\frac{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)u(p_{i})}{\sum_{j:y_{j}\geq y_{i}}\exp\left(x_{j}^{\top}\hat{\beta}+\hat{f}(p_{j})\right)}\right).$ (14)
Define the following quantities:
$\theta_{i}(\beta,f)$ $=\exp(x_{i}^{\top}\beta+f(p_{i})),$
$r_{ji}$ $=I(Y_{j}\geq Y_{i}),$
$s_{i}(\beta,f)$ $=\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f).$
Thus, (13) can be rewritten as:
$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\frac{\delta_{i}}{s_{i}(\beta,f)}\sum_{j=1}^{n}r_{ji}\theta_{j}(\beta,f)x_{jk}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{j=1}^{n}\theta_{j}(\beta,f)x_{jk}\sum_{i=1}^{n}r_{ji}\frac{\delta_{i}}{s_{i}(\beta,f)}$
$=\sum_{i=1}^{n}\delta_{i}x_{ik}-\sum_{i=1}^{n}\theta_{i}(\beta,f)x_{ik}l_{i}(\beta,f)$
$=\sum_{i=1}^{n}(\delta_{i}-\theta_{i}(\beta,f)l_{i}(\beta,f))x_{ik},$
where $l_{i}(\beta,f)=\sum_{j=1}^{n}r_{ij}\frac{\delta_{j}}{s_{j}(\beta,f)}$. Continuing, we have:
$\frac{\partial\mathcal{L}(\beta,f)}{\partial\beta_{k}}$ $=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\theta_{i}(\beta,f)x_{ik}$
$=\sum_{i=1}^{n}\frac{l_{i}(\beta,f)^{-1}\delta_{i}-\theta_{i}(\beta,f)}{\theta_{i}(\beta,f)l_{i}(\beta,f)^{-1}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}.$
Similarly, the Gâteaux derivative of $\mathcal{L}(\beta,f)$ in the direction $u$ is:
$\frac{\partial\mathcal{L}(\beta,u)}{\partial f}=\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}.$
Thus, solving (12) is equivalent to finding $(\hat{\beta},\hat{f})$ that satisfies:
$\begin{cases}\sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,f)}{\partial\beta_{k}}=0,\quad\forall k=1,\dots,q,\\ \sum_{i=1}^{n}\frac{l_{i}^{-1}\delta_{i}-\theta_{i}(\beta,f)}{v_{i}}\frac{\partial\theta_{i}(\beta,u)}{\partial f}+\mu m(u,\hat{f})=0,\quad\forall u\in\mathcal{F}.\end{cases}$ (15)
Assuming $l_{i}$ and $v_{i}$ are constant, solving (15) is equivalent to finding the minimizer of the functional:
$\mathcal{J}_{\mu}(\beta,f)=\|V^{-1/2}(\delta/l-\theta)\|^{2}+\mu m(f,f),$
where $\delta/l=(\delta_{1}/l_{1},\dots,\delta_{n}/l_{n})^{\top}$ and $V=\text{diag}(v_{1},\dots,v_{n})$. Assuming such quantities are constant suggests an iterative computation scheme.
Let $\theta^{(k)}$ be an estimate of $\theta(\beta,f)$ after $k$ iterations. We consider a first-order approximation of $\theta$ near the current value $\theta^{(k)}=(\beta^{(k)},f^{(k)})$:
$\theta(\beta,f)\approx\underbrace{\exp(X\beta^{(k)}+f^{(k)})}_{\theta^{(k)}}+\frac{\partial\theta(\beta,f)}{\partial\beta}(\beta-\beta^{(k)})+\frac{\partial\theta(\beta,f)}{\partial f}(f-f^{(k)}).$
The partial derivatives of $\theta$ with respect to both parameters are computed as follows. For the derivative with respect to $\beta$:
$\frac{\partial\theta_{i}^{(k)}}{\partial\beta_{j}}=\partial\theta_{i}^{(k)}x_{ij},$
which in matrix form is:
$\frac{\partial\theta(\beta,f)}{\partial\beta}=\Theta^{(k)}X,$
where $\Theta^{(k)}=\text{diag}(\theta^{(k)})$.
Similarly, for the derivative with respect to $f$, we have:
$\frac{\partial\theta_{i}(\beta^{(k)},f^{(k)}+\tau(f-f^{(k)}))}{\partial f}=\theta_{i}(\beta^{(k)},f^{(k)})(f(p_{i})-f^{(k)}(p_{i})).$
Hence, the first-order approximation of $\mathcal{J}_{\mu}(\beta,f)$ in the neighborhood of $(\beta^{(k)},f^{(k)})$ is:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|V^{-1/2}\left[\delta/l-\left(\theta^{(k)}+\Theta^{(k)}X(\beta-\beta^{(k)})+\Theta^{(k)}(f-f^{(k)})\right)\right]\right\|^{2}+\mu m(f,f)$
$=\left\|V^{-1/2}\Theta^{(k)}\left((\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}-X\beta-f\right)\right\|^{2}+\mu m(f,f).$
Defining $z^{(k)}=(\Theta^{(k)})^{-1}(\delta/l-\theta^{(k)})+X\beta^{(k)}+f^{(k)}$, and letting ${W}^{(k)}$ be the $n\times n$ diagonal matrix with $i$th entry equal to $v_{i}^{-1}(\beta^{(k)},f^{(k)})\theta_{i}^{2}(\beta^{(k)},f^{(k)})$, we can rewrite:
$\tilde{\mathcal{J}}^{(k)}_{\lambda}(\beta,f)=\left\|({W}^{(k)})^{1/2}(z^{(k)}-X\beta-f)\right\|^{2}+\mu m(f,f).$ (16)
Since ${W}^{(k)}$ is positive definite, $\tilde{\mathcal{J}}_{\lambda}$ is a quadratic form with a unique minimum.